# CDQuant: Accurate Post-training Weight Quantization of Large Pre-trained Models using Greedy Coordinate Descent

### v2

@@ Appendix D Detailed Downstream Evaluation Results
w2a16g128 NatualQ. SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
GPTQ - $11.55$ $66.72$ $48.79$ $7.43$ $47.27$ $76.89$ $84.56$ $74.36$ $79.54$ $73.8$
CD $1$ $11.25$ $64.49$ $50.84$ $7.92$ $48.12$ $75.97$ $79.39$ $74.28$ $79.11$ $73.95$
BCD(k=2) $1$ $11.77$ $64.17$ $51.36$ $8.56$ $48.38$ $75.88$ $81.01$ $74.6$ $78.84$ $74.27$
$2$ $11.61$ $64.69$ $51.23$ $8.51$ $47.87$ $76.09$ $81.74$ $74.72$ $79$ $74.9$
$3$ $11.72$ $64.5$ $52.15$ $8.71$ $48.46$ $76.18$ $82.54$ $74.82$ $79.6$ $74.43$
BCD(k=4) $1$ $11.69$ $64.48$ $51.07$ $8.51$ $47.7$ $75.63$ $80.89$ $74.55$ $79.6$ $74.27$
$2$ $11.75$ $64.39$ $51.36$ $7.97$ $48.46$ $75.04$ $80.58$ $74.48$ $79.38$ $73.88$
$3$ $11.8$ $64.41$ $51.16$ $8.02$ $48.12$ $75.55$ $80.86$ $74.55$ $78.73$ $74.59$
Appendix E Quantization Runtime
Table 19: Table presents the time taken (in minutes) to quantize PaLM2-Otter and PaLM2-Bison. All models were quantized with 8 80GB H100s, and run for 1 epoch (i.e., $T=d_{\text{in}}$).
Config Method PaLM2-Otter PaLM2-Bison
FFN runtime
(in minutes)
Attn. runtime
(in minutes)
FFN runtime
(in minutes)
Attn. runtime
(in minutes)
w3a16 GPTQ $0.90$m $0.72$m $6.42$m $0.63$m
CD $2.94$m $0.49$m $16.13$m $0.43$m
BCD(k=2) $14.03$m $2.08$m $82.13$m $1.81$m
w3a16g128 GPTQ $1.02$m $0.75$m $6.70$m $0.66$m
CD $3.32$m $0.59$m $18.44$m $0.52$m
BCD(k=2) $14.73$m $2.12$m $88.10$m $1.86$m
w4a16 GPTQ $0.90$m $0.64$m $6.60$m $0.56$m
CD $3.59$m $0.59$m $21.19$m $0.52$m
BCD(k=2) $34.85$m $4.72$m $235.14$m $4.13$m
w4a16g128 GPTQ $0.98$m $0.69$m $6.962$m $0.60$m
CD $3.99$m $0.69$m $23.48$m $0.60$m
BCD(k=2) $42.37$m $5.41$m $192.59$m $4.74$m
Appendix F Comparison to QuantEase
We also compare our per-channel quantization results with those of QuantEase, a parallel study to ours. It’s important to note that QuantEase does not have a publicly available implementation. So, we implemented the cyclic coordinate descent strategy used by QuantEase, and used the same initialization and regularization strength as our algorithms (although the QuantEase paper doesn’t provide these details). We then ran QuantEase for the recommended number of iterations specified in the paper ($20$ epochs or $T=20d_{\text{in}}$ iterations). The findings are presented in Table 20. On PaLM2-Gecko, PaLM2-Otter, and PaLM2-Bison, both these approaches have similar performance. These results collectively highlight the potential of coordinate descent algorithms in outperforming GPTQ.
Config Method PaLM2-Gecko PaLM2-Otter PaLM2-Bison
w3a16 GPTQ $11.347$ $7.176$ $5.774$
QuantEase (epochs=20) $10.731$ $6.996$ $5.741$
CD $10.920$ $7.002$ $5.739$
BCD(k=2) $10.898$ $6.979$ $5.733$
w4a16 GPTQ $8.764$ $6.249$ $5.417$
QuantEase (epochs=20) $8.670$ $6.197$ $5.408$
CD $8.694$ $6.195$ $5.407$
BCD(k=2) $8.691$ $6.192$ $5.405$
Table 20: Comparison of QuantEase with CD and BCD for FFN quantization.

### v3

@@ F.2 Detailed Downstream Evaluation Results
NaturalQ SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
w16a16 $26.76$ $72.06$ $76.83$ $21.7$ $58.96$ $77.57$ $83.33$ $77.31$ $81.12$ $67.96$
w316 OWC $21.02$ $74.18$ $69.76$ $18.41$ $56.23$ $76.52$ $84.37$ $75.24$ $80.79$ $67.88$
GPTQ $21.22$ $74$ $69.99$ $18.06$ $57$ $77.19$ $82.35$ $74.96$ $80.69$ $66.46$
CD $22.24$ $73.41$ $70.67$ $19.59$ $56.91$ $77.74$ $84.13$ $75.21$ $79.82$ $66.61$
BCD(k=2) $21.77$ $73.33$ $71$ $18.8$ $57.08$ $77.69$ $83.18$ $75.04$ $80.03$ $67.01$
w3a16g128 OWC $23.1$ $73.42$ $71.79$ $19.29$ $56.31$ $76.6$ $83.94$ $75.28$ $80.74$ $67.25$
GPTQ $22.33$ $72.77$ $71.68$ $19.69$ $56.66$ $77.44$ $83.46$ $75.19$ $80.74$ $67.64$
CD $22.8$ $71.94$ $72.24$ $19.39$ $56.74$ $77.36$ $84.13$ $75.27$ $80.14$ $66.38$
BCD(k=2) $23.02$ $72.3$ $71.99$ $19.29$ $56.57$ $77.06$ $84.37$ $75.36$ $80.3$ $66.46$
OWC-CD $22.55$ $72.69$ $72.13$ $19.19$ $55.46$ $76.85$ $82.81$ $75.36$ $80.79$ $66.93$
OWC-CD + CD $22.99$ $72.42$ $72.28$ $19.09$ $56.91$ $77.36$ $82.75$ $75.48$ $80.52$ $67.72$
OWC-CD + BCD(k=2) $23.27$ $72.39$ $72.23$ $18.9$ $57$ $77.31$ $83.27$ $75.49$ $80.41$ $67.32$
w4a16 OWC $25.48$ $73.21$ $74.93$ $20.32$ $58.45$ $76.81$ $83.94$ $76.58$ $80.69$ $68.11$
GPTQ $25.29$ $72.9$ $74.65$ $21.16$ $58.7$ $77.27$ $83.55$ $76.6$ $81.23$ $68.98$
CD $25.57$ $73.03$ $75.42$ $20.92$ $58.36$ $77.61$ $83.79$ $76.75$ $80.41$ $68.67$
BCD(k=2) $25.6$ $72.86$ $75.35$ $21.06$ $58.62$ $77.4$ $83.73$ $76.74$ $80.69$ $68.59$
w4a16g128 OWC $26.26$ $73.33$ $75.73$ $21.56$ $58.7$ $77.4$ $83.33$ $76.76$ $80.85$ $66.85$
GPTQ $26.12$ $72.53$ $75.75$ $21.21$ $58.53$ $77.65$ $83.06$ $76.72$ $81.28$ $66.61$
CD $26.23$ $72.57$ $75.83$ $20.96$ $58.28$ $77.53$ $83.15$ $76.83$ $81.12$ $67.4$
BCD(k=2) $26.4$ $72.62$ $75.94$ $21.11$ $58.36$ $77.4$ $83.09$ $76.72$ $80.85$ $67.48$
OWC-CD $26.15$ $73.46$ $75.88$ $21.56$ $59.3$ $77.74$ $82.81$ $76.72$ $80.96$ $67.72$
OWC-CD + CD $26.15$ $72.7$ $75.98$ $21.41$ $58.62$ $77.78$ $82.94$ $76.86$ $81.07$ $67.48$
OWC-CD + BCD(k=2) $26.32$ $72.75$ $76.13$ $21.21$ $59.13$ $77.48$ $83$ $76.92$ $81.01$ $67.48$
Table 16: Table presents downstream evaluation (zero-shot) numbers of Gemma-2 27B - for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights.
Config Method Gemma-2 27B
NaturalQ SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
w16a16 $30.42$ $75.04$ $81.5$ $23.33$ $60.49$ $78.96$ $83.18$ $82.24$ $84.6$ $75.69$
w316 OWC $23.3$ $73.28$ $73.2$ $19.39$ $56.66$ $77.53$ $81.01$ $79.82$ $83.24$ $73.01$
GPTQ $26.09$ $75.08$ $76.62$ $21.85$ $57.51$ $78.96$ $77.4$ $80.57$ $83.73$ $74.98$
CD $26.51$ $75.57$ $77.21$ $21.06$ $59.13$ $79.76$ $78.29$ $80.1$ $83.95$ $74.66$
BCD(k=2) $26.7$ $75.67$ $77.34$ $21.65$ $58.45$ $79.34$ $78.47$ $79.93$ $83.41$ $76.56$
w3a16g128 OWC $26.7$ $75.95$ $78.61$ $21.8$ $58.45$ $79.63$ $83.55$ $80.49$ $82.75$ $76.01$
GPTQ $27.48$ $75.97$ $78.72$ $22.88$ $59.56$ $80.05$ $80.28$ $81.03$ $83.68$ $74.35$
CD $27.98$ $76.15$ $78.59$ $22.54$ $59.73$ $79.88$ $81.77$ $80.99$ $84.11$ $76.09$
BCD(k=2) $27.78$ $75.99$ $78.62$ $22.49$ $59.56$ $80.01$ $81.99$ $81.08$ $84.11$ $76.56$
OWC-CD $26.87$ $76.03$ $79$ $22.05$ $58.96$ $80.18$ $81.28$ $80.94$ $83.79$ $76.64$
OWC-CD + CD $27.87$ $76.36$ $79.62$ $22.34$ $59.3$ $80.05$ $81.59$ $80.97$ $83.79$ $76.72$
OWC-CD + BCD(k=2) $27.7$ $76.45$ $79.38$ $22.44$ $59.73$ $80.05$ $81.83$ $80.98$ $83.57$ $76.56$
w4a16 OWC $29$ $75.51$ $80.05$ $21.95$ $59.39$ $79.04$ $81.13$ $81.93$ $84.39$ $75.53$
GPTQ $29.5$ $75.7$ $80.41$ $22.88$ $59.47$ $78.7$ $81.62$ $81.91$ $84.33$ $75.85$
CD $29.36$ $75.23$ $80.23$ $22.59$ $59.9$ $79.67$ $81.53$ $81.84$ $84.49$ $75.77$
BCD(k=2) $29.42$ $75.34$ $80.31$ $22.54$ $59.3$ $79.46$ $81.83$ $81.72$ $84.44$ $75.93$
w4a16g128 OWC $29.25$ $75.41$ $80.91$ $22.64$ $59.98$ $79.46$ $83.76$ $82.14$ $84.22$ $76.01$
GPTQ $29.53$ $75.61$ $80.85$ $22.49$ $60.41$ $79.59$ $83.61$ $82.14$ $84.66$ $75.37$
CD $29.58$ $75.38$ $81.05$ $23.13$ $59.98$ $79.42$ $83.73$ $82.08$ $84.44$ $75.93$
BCD(k=2) $29.61$ $75.44$ $80.91$ $22.93$ $60.32$ $79.71$ $83.85$ $82.06$ $84.44$ $75.53$
OWC-CD $29.11$ $75.93$ $80.85$ $22.69$ $60.75$ $79.67$ $83.33$ $82.2$ $84.87$ $75.69$
OWC-CD + CD $29.36$ $75.66$ $80.96$ $22.98$ $59.9$ $79.59$ $83.82$ $82.05$ $85.15$ $75.77$
OWC-CD + BCD(k=2) $29.34$ $75.41$ $80.86$ $22.93$ $59.9$ $79.84$ $83.94$ $82.14$ $84.93$ $75.37$
F.3 Attention + FFN Quantization
Before presenting the results for quantization of both attention and FFN layers, we first explain why quantizing the attention weights does not yield inference latency gains.
F.3.1 Attention Quantization doesn’t yield Latency Gains
In Table 17, we present the end-to-end prefix and decode latency for Gemma-2 27B with three different prefix lengths. The latencies are obtained by running Gemma-2 27B on 4 TPUv5 chips with a per TPU batch size of 4 and no model parallelism.
Prefix phase. During the prefix phase, neither FFN nor attention quantization yielded significant latency improvements, likely due to the compute-bound nature of the prefix.
Decode phase. In the decode phase, quantizing attention weights from 16 bits to 8 bits provided marginal gains, as attention compute is primarily memory-bound by the KV cache size, which typically exceeds attention weight size for moderate-to-large context lengths. Further reduction to 4-bit quantization offered negligible latency improvement while incurring a non-trivial drop in quality (Table 18 and Table 2), thus questioning the utility of attention weight quantization. Conversely, FFN layers, processing only a single token activation during decoding, are memory-bound by the significantly larger FFN weights (9x the size of attention weights in Gemma-2 27B). Consequently, FFN weight quantization effectively reduces memory transfers and improves latency. We observed a $32\%-35\%$ latency reduction when quantizing from 16 bits to 8 bits, with an additional $10\%$ reduction when using 4-bit quantization.
These results clearly highlight that quantization of attention weights doesn’t yield any inference latency gains.
Table 17: Table presents the prefix and decode time inference latency of Gemma-2 27B with three different prefix lengths. In the Attention-only setting, only the attention parameters are quantized and rest of the model is in 16 bits. Similarly, in the FFN-only setting, only the FFN parameters are quantized.
Quantization Prefix Config Prefix Decode
Setting Length time (ms) per step time (ms)
Attention-only 1024 w16a16 $299.958$ $26.178$
w8a16 $302.199$ $25.710$
w4a16 $301.899$ $25.595$
2048 w16a16 $616.460$ $26.532$
w8a16 $616.080$ $26.117$
w4a16 $615.601$ $25.964$
4096 w16a16 $1296.000$ $27.254$
w8a16 $1290.000$ $26.685$
w4a16 $1286.000$ $26.653$
FFN-only 1024 w16a16 $299.958$ $26.178$
w8a16 $292.630$ $17.410$
w4a16 $284.680$ $15.635$
2048 w16a16 $616.460$ $26.532$
w8a16 $604.941$ $17.868$
w4a16 $602.306$ $16.006$
4096 w16a16 $1296.000$ $27.254$
w8a16 $1276.000$ $18.521$
w4a16 $1273.000$ $16.626$
F.3.2 Results for Attention + FFN quantization
We now perform FFN + Attention weight quantization, results for which are presented in Table 18. We use the same experimental setting as the one described in Section 4. In our experiments we noticed that the inputs to the attention layer are often aligned in a handful of directions. Consequently, performing quantization using such a data leads to a huge drop in performance, as the algorithms would primarily focus on a few directions and ignore the rest. To mitigate this, we clip the largest eigenvalues of the Hessian to ensure a more balanced Hessian. This technique, reminiscent of the weight clipping in OmniQuant (Shao et al., 2023), improves the performance of both GPTQ and CDQuant [^1] One could also rely on existing techniques such as AWQ and SmoothQuant to reduce the effect of outliers. But in our experiments we noticed that both AWQ and SmoothQuant performed poorly compared to the simple eigenvalue clipping technique.. For instance, for PaLM2-Gecko, the perplexity for w3a16-GPTQ improves from $34.872$ to $24.054$ with clipping, and for w4a16-GPTQ, it improves from $10.966$ to $10.005$. Table 18 presents the results from this experiment, where we run all the algorithms on the clipped Hessian. We once again notice that in almost all the settings, our coordinate descent algorithms outperform GPTQ. For instance, we see $5-10\%$ improvement in perplexity for 3-bit per-channel quantization of Gemma-2, and PaLM2 Gecko models.
Table 18: Table presents the perplexity evaluations for GPTQ, CD, BCD for INT3, INT4 quantization of both FFN and attention weights.
Config Method Gemma-1 Gemma-2 Gemma-2 PaLM2 PaLM2 PaLM2
7B 9B 27B Gecko Otter Bison
w16a16 - $10.348$ $10.683$ $8.682$ $7.948$ $5.984$ $5.298$
w3a16 OWC $61.89e$ 6 $12.558$ $15.736$ $41.751$ $54.499$ $6.446$
GPTQ $91.586$ $11.849$ $11.423$ $24.054$ $\mathbf{11.384}$ $5.801$
CD $36.102$ $11.808$ $10.941$ $19.692$ $11.542$ $5.770$
BCD(k=2) $\mathbf{32.850}$ $\mathbf{11.784}$ $\mathbf{10.926}$ $\mathbf{19.312}$ $11.392$ $\mathbf{5.762}$
w3a16g128 OWC $33.710$ $11.792$ $10.486$ $20.301$ $9.785$ $5.973$
GPTQ $20.045$ $11.491$ $9.766$ $15.042$ $7.501$ $5.698$
CD $15.850$ $11.500$ $9.689$ $14.757$ $7.454$ $5.679$
BCD(k=2) $15.466$ $11.487$ $9.657$ $14.691$ $\mathbf{7.438}$ $5.673$
OWC-CD $18.599$ $11.583$ $9.820$ $16.268$ $7.617$ $5.746$
OWC-CD + CD $15.792$ $11.465$ $\mathbf{9.559}$ $14.285$ $7.466$ $5.675$
OWC-CD + BCD(k=2) $\mathbf{15.777}$ $\mathbf{11.463}$ $9.564$ $\mathbf{14.242}$ $7.458$ $\mathbf{5.670}$
w4a16 OWC $40.563$ $11.100$ $9.706$ $10.683$ $7.194$ $5.508$
GPTQ $15.805$ $11.021$ $9.217$ $10.005$ $6.645$ $5.425$
CD $14.379$ $\mathbf{10.979}$ $9.177$ $9.991$ $6.627$ $5.415$
BCD(k=2) $\mathbf{14.056}$ $10.982$ $\mathbf{9.167}$ $\mathbf{9.965}$ $\mathbf{6.621}$ $\mathbf{5.413}$
w4a16g128 OWC $12.222$ $10.919$ $9.027$ $9.402$ $6.415$ $5.420$
GPTQ $11.548$ $10.863$ $8.888$ $9.160$ $6.241$ $5.381$
CD $11.306$ $10.856$ $8.890$ $9.124$ $6.226$ $5.377$
BCD(k=2) $11.105$ $\mathbf{10.853}$ $8.892$ $9.120$ $\mathbf{6.225}$ $5.376$
OWC-CD $11.342$ $10.876$ $8.924$ $9.241$ $6.261$ $5.387$
OWC-CD + CD $11.149$ $10.863$ $8.873$ $9.094$ $6.244$ $5.375$
OWC-CD + BCD(k=2) $\mathbf{11.140}$ $10.866$ $\mathbf{8.873}$ $\mathbf{9.093}$ $6.242$ $\mathbf{5.374}$
F.4 Runtime Analysis
In Tables 19 20, we present the quantization runtime of GPTQ, CD and BCD on Gemma models using H100s (in one setting we used 1 H100 and in another we used 8H100s). In particular, in Table 19, we present runtime for quantizing all the attention and FFN layers (which includes FFN1, FFN1-gate, FFN2) in the model. Table 20 presents a more detailed breakdown; we present the per-layer runtime for quantizing FFN1 and FFN2 layers.
Runtime on $8$ H100’s. With 8 H100s, CD is as fast as GPTQ for quantizing attention weights, whereas BCD is 2×2$\times$2 × slower. However, for FFN1 (FFN2) quantization, CD is 5×5$\times$5 × (2×2$\times$2 ×) slower than GPTQ, whereas BCD is an order of magnitude slower than GPTQ. For quantization of the entire model, CD is 2×2$\times$2 × slower than GPTQ (because quantizing FFN2 takes up most of the time). We attribute the slowness of CD and BCD to the numerous scatters and gathers involved in finding the most optimal coodrinate/block of coordinate to update, and then eventually updating it. Since ML accelerators are slow with gather and scatter operations, we are bound by the time it takes to do them. Performing these operations on the CPU should help speed up CD and BCD.
Runtime on $1$ H100. With a single H100, GPTQ is $2$x-$3.8$x faster than CD for the smaller, attention weights, and is an order of magnitude faster than BCD. For the larger, FFN weights, GPTQ is an order of magnitude faster than CD, and in most cases is two orders of magnitude faster than BCD. As stated previously, both CD and BCD are bound by the time it takes to do gathers and scatters.
Speeding up BCD. To algorithmically speed up BCD, we propose to use CD for FFN2 quantization, and use BCD for the rest of the layers. We find that BCD spends most of its time quantizing the FFN2 weight matrix (as can be seen in Table 20) where the quantization axis is of size $d_{\text{hidden}}$. In most cases, the time it takes to FFN2 is $10$x the time it takes to quantize FFN1. Thus, to speed up BCD, we quantize FFN2 with CD, and FFN1 and FFN1-gate with BCD (where the quantization axis is of size $d_{\text{model}}$. Results for this setting can be found in Table 21. It can be seen that quantizing FFN2 with CD leads to negligible to no drop in performance especially for the larger, and better Gemma-2 9B. In fact, it improves the performance of BCD for w3a16g128. With this setup, BCD’s runtime is reduced significantly and is comparable to that of CD.
Speeding up CD. We find that CD converges to an optimal solution significantly before $d_{\text{in}}$ iterations. Based on this finding, we run CD for lesser number of iterations. In Table 22, one can see that there is negligible quality drop for CD, and even with $d_{\text{in}}/8$ iterations, CD outperforms GPTQ. With lesser number of iterations, CD’s runtime can be made faster than/comparable to that of GPTQ.
Table 19: Table presents the per-channel quantization runtime (in minutes) of GPTQ, CD, BCD (where each run is for $d_{\text{in}}$ iterations). FFN includes the combined runtime of quantizing FFN1, FFN1-gate and FFN2 layers.
Config Method 8 H100s 1 H100
Gemma-1 7B Gemma-2 9B Gemma-2 27B Gemma-1 7B Gemma-2 9B Gemma-2 27B
FFN
Attn.
FFN
Attn.
FFN
Attn.
FFN
Attn.
FFN
Attn.
FFN
Attn.
w3a16 GPTQ $1.68$ $0.67$ $1.15$ $1.05$ $7.98$ $1.18$ $2.13$ $0.65$ $1.36$ $1.04$ $10.31$ $1.32$
CD $3.03$ $0.19$ $2$ $0.31$ $16.55$ $0.51$ $22.3$ $1.2$ $13.31$ $1.87$ $122.35$ $3.2$
BCD(k=2) $8.52$ $0.71$ $5.95$ $1.18$ $66.97$ $2.11$ $88.14$ $3.85$ $61.5$ $7.24$ $541.03$ $13.62$
w4a16 GPTQ $1.69$ $0.77$ $1.15$ $1.06$ $7.95$ $1.18$ $2.1$ $0.63$ $1.36$ $0.98$ $10.16$ $1.12$
CD $3.9$ $0.22$ $2.53$ $0.36$ $21.16$ $0.64$ $29.27$ $1.47$ $17.66$ $2.25$ $163.78$ $4.27$
BCD(k=2) $26.2$ $1.76$ $18.51$ $2.66$ $162.43$ $4.64$ $218.25$ $12.89$ $163.64$ $20.12$ $1296.36$ $36.68$
Table 20: Table presents the per-layer, per-channel quantization runtime (in seconds) of GPTQ, CD, BCD (where each run is for $d_{\text{in}}$ iterations) for the FFN layer. Note that FFN1 and FFN1-Gate have the exact same runtimes for all compared methods, hence we report only FFN1’s runtime.
Config Method 8 H100s 1 H100
Gemma-1 7B Gemma-2 9B Gemma-2 27B Gemma-1 7B Gemma-2 9B Gemma-2 27B
FFN1
FFN2
FFN1
FFN2
FFN1
FFN2
FFN1
FFN2.
FFN1
FFN2
FFN1
FFN2
w3a16 GPTQ $0.34$ $2.93$ $0.35$ $0.94$ $0.39$ $9.61$ $0.43$ $3.69$ $0.39$ $1.16$ $0.81$ $11.64$
CD $0.46$ $5.57$ $0.41$ $2.04$ $1.6$ $18.38$ $3.34$ $41.11$ $2.86$ $13.31$ $12.24$ $135.11$
BCD(k=2) $1.4$ $15.46$ $1.27$ $5.96$ $7.86$ $71.67$ $12.55$ $163.74$ $11.95$ $63.95$ $57.38$ $590.89$
w4a16 GPTQ $0.35$ $2.93$ $0.36$ $0.94$ $0.4$ $9.58$ $0.42$ $3.68$ $0.39$ $1.14$ $0.81$ $11.63$
CD $0.53$ $7.28$ $0.46$ $2.71$ $2.15$ $23.31$ $3.94$ $54.85$ $3.22$ $18.81$ $17.24$ $179.15$
BCD(k=2) $4.54$ $47.07$ $3.71$ $19.02$ $18.97$ $173.92$ $35.95$ $395.78$ $30.34$ $173.09$ $151.36$ $1388.19$
Table 21: Table presents the perplexity evaluations for INT3 quantization with BCD wherein the FFN2 may be quantized either with CD or BCD, while rest of the FFN parameters are quantized with BCD.
Config Method Gemma-1 7B Gemma-2 9B
w16a16 $10.384$ $10.683$
w3a16 CD, FFN2 $19.709$ $11.288$
BCD(k=2), FFN2 $18.552$ $11.284$
w3a16g128 CD, FFN2 $13.512$ $11.179$
BCD(k=2), FFN2 $13.501$ $11.180$
Table 22: Table presents the perplexity evaluations for INT3 quantization with CD, wherein instead of running CD for $d_{\text{in}}$ iterations, we run it for $d_{\text{in}}/2$, $d_{\text{in}}/4$, and $d_{\text{in}}/8$ iterations.
Config Method Epochs Gemma-1 7B Gemma-2 9B
w16a16 $10.384$ $10.683$
w3a16 GPTQ - $48.157$ $11.382$
CD $1$ $19.614$ $11.301$
$1/2$ $19.614$ $11.301$
$1/4$ $19.621$ $11.301$
$1/8$ $19.597$ $11.305$
w3a16g128 GPTQ - $15.561$ $11.193$
CD $1$ $13.496$ $11.182$
$1/2$ $13.488$ $11.189$
$1/4$ $13.480$ $11.189$
$1/8$ $13.560$ $11.188$